# Disentangling Rich Dynamics from Feature Learning: A Framework for Independent Measurements

### v2

@@ 5 Observations from our visualization methods
Figure 6: VGG16 on CIFAR-100 with and without batch normalization. Adding batch normalization results in not only a significant performance boost, but shifts the model from the lazy to rich regime ($\mathcal{D}_{LR}$ drops from 0.66 to 0.07). The visualization highlights this shift, showing that only a minimal number (100) of features are actively used when batch normalization is present.
@@ 6 Discussion and Conclusion
In this paper, we introduced a principled framework for measuring dynamical richness independently of representation improvement. Our low-rank-based metric $\mathcal{D}_{LR}$, along with a complementary function-space visualization, provides tools to analyze training dynamics and feature improvement separately within a unified framework.
Our metric builds on the low-rank bias in rich regimes and reduces to neural collapse under mild assumptions (Propositions 1 and 2), offering generalization of priors studies. Unlike prior proxies — e.g., parameter norms or neural collapse metrics — $\mathcal{D}_{LR}$ operates directly in function space, assessing how efficiently the model expresses its learned function using minimal number of features.
Empirically, our richness metric decouples performance from dynamics (Figs. 1 and 3), captures lazy-to-rich transitions (Figs. 2 and 4), and detects richness where prior measures fail (Fig. 2). Our visualizations highlight these differences more directly, revealing novel patterns (Figs. 6, 5 and 7) that invite further theoretical study of feature learning.
Limitations.
The current form of $\mathcal{T}_{MP}$ is limited to orthogonal and isotropic target functions. While this covers most classification tasks, a more general setup would be preferable. Additionally, our framework provides a lightweight metric but focuses only on last-layer features, leaving the dynamics of the entire network dynamics unexplored.
@@ Conclusion.
Our metric and visualizations offer a concrete step toward understanding feature learning from both representational and dynamical perspectives. We believe these tools can help bridge empirical findings with theoretical insights and inspire new discoveries through the details they reveal. In future work, we will address current limited scope of balanced tasks.
@@ Acknowledgment
We thank Charles London, Zohar Ringel, and Shuofeng Zhang for their helpful comments. NF acknowledges the UKRI support through the Horizon Europe guarantee Marie Skłodowska-Curie grant (EP/X036820/1). SL was supported by the National Research Foundation of Korea (NRF) grants No.2020R1A5A1016126 and No.RS–2024-00462910. The authors would also acknowledge support from His Majesty’s Government in the development of this research. CM acknowledges funding from an EPSRC iCASE grant with IBM (grant number EP/S513842/1). SH gratefully acknowledge partial support from NSF grants DMS-2209975, 2015341, 20241842, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, and NSF grant MC2378 to the Institute for Artificial CyberThreat Intelligence and OperatioN (ACTION). The authors would also acknowledge support from His Majesty’s Government in the development of this research.
@@ Appendix A Glossary
$C$ Class count The number of classes. Section 2
$p$ Layer width The width of the last layer Section 2
$n$ Sample count The number of training samples Section 2
$\mathcal{X}$ Input space Space of inputs Section 2
$q$ Input distribution The probability distribution that generates samples in the input space. Section 2
$\Phi$ Feature map A map from input to post-activation of the penultimate layer (the activations fed to the last layer). Section 2
@@
A $p$-dimensional random variable or the post-activation of penultimate layer for $x\sim q$. Section 2
$\mathcal{T}$ (Feature kernel) integral operator A map from Hilbert space to Hilbert space that depends on $\Phi(x),~{}x\sim q$ (typically the last layer features) Eq. 1
$\mathcal{T}_{MP}$ Minimum projection operator A special set of operators that depends on the learned function $\hat{f}$. Definition 1
$e_{k}$ $k^{\mathrm{th}}$ eigenfunction The eigenfunction of $\mathcal{T}.$ Eq. 2
$\rho_{k}$ $k^{\mathrm{th}}$ eigenvalue The eigenvalue of $\mathcal{T}$. Eq. 2
$\hat{f}$ Learned function A $C$-dimensional vector output function $\hat{f}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ expressed by the neural network. Section 2
$f^{*}$ Target function A $C$-dimensional vector output function $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ with correct labels. The output is always a one-hot vector (up to a scaling constant). Section 2
$\mathcal{H}^{*}$ Target function space Space linearly spanned by the entries of the target function. Section 3
$\hat{\mathcal{H}}$ Learned function space Space linearly spanned by the entries of the learned function Section 3
$P_{\mathcal{H}}$ Projection operator Projection operator onto $\mathcal{H}$. If $\mathcal{H}=span\{e_{1},e_{2},\ldots,e_{p}\}$, where $e_{k}$’s are orthonormal, the the projection operator is given as $P_{\mathcal{H}}=\sum_{k=1}^{p}\ket{e_{k}}\bra{e_{k}}$ |. Section 3
$\alpha$ Downscale constant Prefactor that downscaling the target such that $y\rightarrow y/\alpha$ Section 4
$\Sigma_{b}$ Inter-class covariance matrix Covariance of class mean vectors for the training set. Eq. 34
$\Sigma_{W}$ Intra-class covariance matrix Covariance of training feature vectors within given class. Eq. 33
@@
$\|\theta\|_{F}^{2}$ Parameter norm Norm of all parameters in the model. It was used as richness measure with smaller value meaning richer dynamics. Section 4
$\mathcal{D}_{LR}$ low rank metric Our proposed metric of dynamical richness. Smaller is richer. Eq. 4
$\mathcal{S}_{init}$ Kernel deviation The CKA measure between the learned and initial kernel. Section 4
$NC1$ Neural collapse metric The trace of $\Sigma_{b}^{\dagger}\Sigma_{W}$, measuring the training feature vectors variance compared to class boundaries. Proposition 1
$\Pi^{*}(k)$ Cumulative quality Measure of how well the first $k$ eigenfunctions span the target function space. Eq. 5
$\hat{\Pi}(k)$ Cumulative utilization Measure of how well the first $k$ eigenfunctions span the learned function space. Eq. 5
$Q^{*}(k)$ Quality Per-feature quality or $\Pi^{*}(k)-\Pi^{*}(k-1)$ where $\Pi^{*}(0)=0$ Fig. 7
$\hat{Q}(k)$ Utilization Per-feature utilization or $\hat{\Pi}(k)-\hat{\Pi}(k-1)$ where $\hat{\Pi}(0)=0$ Fig. 7
@@ B.1 Bra-ket notation
$\mathcal{T}[f](x)=[\Phi_{1}(x)\braket{\Phi_{1}}{f},\ldots,\Phi_{p}(x)\braket{\Phi_{p}}{f}],$
where $\braket{\Phi_{k}}{f}$ are scalars and $\Phi_{k}:\mathcal{X}\rightarrow\mathbb{R}$ are functions.
@@ B.2 Features are random variables and (well-behaved) random variables form a Hilbert space
For an underlying probability distribution and a set of functions (feature maps), we can form a Hilbert space. A common example of a Hilbert space will be the set of solutions expressed by a linear model such as
$f(x)=w_{1}x+w_{2}x^{2}+w_{3}x^{3}.$
Assuming the features $[x,x^{2},x^{3}]$ are linearly independent on input distribution $q$, they form a 3-dimensional Hilbert space.
B.3 Centered kernel alignment
The centered kernel alignment ($CKA$[36] measures the similarity between two matrices or operators, and are commonly used for tracking the evolution of NTK [50, 51]:
$CKA(A,B)=\frac{\mathrm{Tr}(c(A)c(B))}{\|c(A)\|_{F}\|c(B)\|_{F}},$ (7)
where $c(A)=(I-\ket{\mathbf{1}}\bra{\mathbf{1}})A(I-\ket{\mathbf{1}}\bra{\mathbf{1}})$ is centering operator, where $I$ is identity, |$\ket{\mathbf{1}}$ is the constant function or constant vector, and $\|\cdot\|_{F}$ is Frobenius norm. The centering operator removes the constant shift. The analogy is measuring the similarity of two multivariate Gaussian random variables by comparing their covariance matrices $\mathbf{E}[(X-\mathbf{E}[X])(X-\mathbf{E}[X])^{T}]$ instead of their autocorrelation matrices $\mathbf{E}[XX^{T}]$.
For $\mathcal{T}_{MP}$, if the learned function space $\hat{\mathcal{H}}$ contains the constant function — which is empirically true throughout the training for all our experiments and always true if it perfectly fits the training samples (Appendix D) — $c(\mathcal{T}_{MP})/\|c(\mathcal{T}_{MP})\|_{F}$ becomes an orthogonal projection operator on the constant complement of $\hat{\mathcal{H}}$, and is independent of the values of $a_{1}$ and $a_{2}$.
@@ B.4 Effective dimension
A covariance matrix or an integral operator may be fully ranked, but their eigenvalues may decay fast (e.g., by a power-law). The small eigendirections are often negligible, and researchers use various effective dimensions to measure the number of significant dimensions. In this appendix, we will use the exponent of entropy [52] to measure the effective dimension.
For a matrix or an operator with positive eigenvalues $[\rho_{1},\cdots,\rho_{p}]$, the effective dimension is
@@
The effective dimension of a matrix (operator) can be interpreted as the number of linearly independent vectors (functions) needed to effectively describe the matrix (operator). This is similar in spirit to the principal component analysis (PCA) [53] in that only the directions with significant variance are considered.
As discussed in Appendix D, the first eigenfunction is always a constant function, which is irrelevant for our measure $\mathcal{D}_{LR}$ and dynamical interpretation. Because the first eigenvalue is often much larger than other significant eigenvalues, we use
@@ Appendix C Linear models, features, and their inductive bias
Here, we will introduce the recent findings on the inductive bias of linear models and demonstrate the significance of eigenvalues and eigenfunctions of $\mathcal{T}$. Linear regression or kernel regression is a rare case where its inductive bias is analytically calculable [16, 17, 18, 19, 20, 21, 54]. For a given feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$ or kernel $K(x,x^{\prime})=\Phi(x)^{T}\Phi(x^{\prime})$, the model is expressed as the $p$ dimensional ($p$ can be infitnite) linear model
@@ C.1 Eigenfunctions - orthonormalized features
In many cases, $[\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)]$ are not orthonormal. For example, the features $[x,x^{2}]$ of $f(x)=w_{1}x+w_{2}x^{2}$ for $q=unif[0,1]$ are not orthonormal:
[…]
The diagonalization into the eigenfunctions returns the orthonormal basis of the hypothesis (Hilbert) space $\mathcal{H}=span\{\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)\}$, more formally known as Reproducing Kernel Hilbert Space (RKHS). The eigenvalue now shows the norm of the feature along the direction of eigenfunction as the features $[\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)]$ are not normalized (Fig. 8).
Figure 8: Diagonalization of features Since the features $[\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)]$ span a vector space, each function $\Phi_{k}:\mathcal{X}\rightarrow\mathbb{R}$ can be represented as a vector. The overlap among $\Phi_{k}$’s (non-zero linear correlation) and differing norms of $\Phi_{k}$’s result in diagonalized features $[\sqrt{\rho_{1}}e_{1},\sqrt{\rho_{2}}e_{2},\cdots,\sqrt{\rho_{p}}e_{p}]$ to have distinct intensities (norms) $\rho_{1},\cdots,\rho_{p}$.
@@
$\Phi_{i}(x)=\sum_{j}O_{ij}\sqrt{\rho_{j}}e_{j}(x),\quad\quad O_{ij}=\frac{1}{\sqrt{\rho_{j}}}\braket{\Phi_{i}}{e_{j}},$ (13)
where $O\in\mathbb{R}^{p\times p}$ is an orthogonal matrix (follows trivially from Eq. 2). Using Eq. 13, the model can be reparameterized in the eigenfunction basis
@@
Note that because $\theta=Ow$, the norm of the parameters is conserved (i.e. $\|\theta\|_{F}=\|w\|_{F}$), indicating that weighted basis $[\sqrt{\rho_{1}}e_{1},\sqrt{\rho_{2}}e_{2},\cdots,\sqrt{\rho_{p}}e_{p}]$ instead of normalized basis $[e_{1},e_{2},\cdots,e_{p}]$ must be used to represent the ‘intensity’ or norm of the features $[\Phi_{1}(x),\cdots,\Phi_{p}(x)]$.

### v3

@@ 5.3 Visualization results
Figure 3: Visualization of VGG16 on CIFAR-100 with and without batch normalization. We visualize the last row of Table˜3, where batch normalization shifts the model from the lazy to the rich regime. The eigenvalue distribution (iii) highlights this difference: with batch normalization, only 100 features are significant, whereas without it the eigenvalues decay slowly.
Figure 4: Visualization on the role of learning rate. We visualize the $2^{nd}$ row of Table˜3 where the learning rates are varied (up to training instability) for ResNet18 on CIFAR-100. The second column (ii) shows that smallest learning rate uses significantly more eigenfunctions (features), while other models uses minimal 100 eigenfunctions, indicating a lazier dynamics.
@@
Figs.˜4 and 3 illustrate how visualization complements the metric. In Fig.˜3, the batch-norm included model uses only 100 features (eigenfunctions) in (ii) and shows clear low-rank structure in (iii). The batch normalization lacking model uses significantly more feature (4096 in total) in (ii) and show power-law distribution in (iii): indicating significantly lazy dynamics. In Fig.˜4(ii), the model with the smallest learning rate uses over 100 features to express the learned function, unlike models with larger learning rates, suggesting lazier dynamics and poorer performance.
Fig.˜5 reveals another novel pattern in feature learning dynamics: feature quality correlates with feature intensity during training, with larger features improving faster. While it is expected that a generalizing model in the rich regime obtains a few high-quality features after training, the correlation between quality and intensity during training has not been previously observed or studied. Interestingly, it differs from the silent alignment effect (Atanasov et al., 2021), where, in the dynamics of linear networks in a rich regime, quality precedes intensity.
@@ 6 Discussion and Conclusion
We introduced a performance-independent metric for dynamical richness, $\mathcal{D}_{LR}$, which reduces to neural collapse as a special case (Propositions˜1 and 2) and empirically agrees with prior studies on lazy-to-rich transitions (Tables˜2 and 3).
Table˜3 and Fig.˜2 provide examples of how the metric can be used for the analysis. We confirm established assumptions explicitly with an independent metric, while potentially challenging a few (Fig.˜5). We also reveal new findings (Fig.˜3), offering a direction for theoretical research.
Moreover, we provide a complementary visualization to extend these results for additional information, also uncovering novel insights such as the alignment between representation quality and feature intensity (Fig.˜5).
Interpretation of neural collapse.
We share our secondary findings on neural collapse. The functional utility of neural collapse remains an open question, particularly regarding its debated link to generalization (Hui et al., 2022; Kothapalli et al., 2022; Zhu et al., 2021). As established in Propositions˜1 and 2, our metric is mathematically tied to neural collapse. Because our metric quantifies representational richness, this suggests neural collapse may serve fundamentally as an indicator of rich feature dynamics rather than generalization.
Intermediate layers.
While our metric achieves computational efficiency by focusing exclusively on the last layer, this design inherently omits the dynamics of intermediate layers. In Appendix˜I, we share our exploration of the intermediate layers and the justification of our choice. Theoretically, our kernel-based framework assumes linearity between features and the learned function — a premise less applicable to intermediate layers subject to non-linear transformations — and empirically, intermediate layers yield negligible additional insight into dynamical richness. We attribute this to the last layer exhibiting the strongest low-rank bias, a view supported by the comparative rarity of deep neural collapse Rangamani et al. (2023) and the shallow nature of unlearning Jung et al. (2025). We leave the extension of this framework to deep feature learning dynamics for future work.
Isotropic target function.
The current form of $\mathcal{T}_{MP}$ is limited to orthogonal and isotropic target functions. While this covers most classification tasks, a more general setup would be preferable. However, imbalanced label remains an active area of research in neural collapse (Fang et al., 2021; Yang et al., 2022; Hong and Ling, 2023; Yan et al., 2024; Gao et al., 2024), and we leave the extension for future work. See Section˜D.4 for further discussion.
@@ Conclusion.
By offering a lightweight, robust, and performance-independent metric on rich dynamics, we aim to lay the groundwork for future theoretical studies on their connection to representation learning. More broadly, we see this work as a diagnostic tool toward bridging empirical observations on representation learning (e.g., rule of thumb) and theoretical understanding of rich dynamics. In future work, we plan to extend beyond the current focus on balanced tasks.
Reproducibility statement
As mentioned in the main text, full experimental details, link to the source code, and statistical significance for the tables are provided in Appendix˜H. All error bars represent one standard deviation. The algorithm for approximating the metric and visualization is provided in Appendix˜E.
@@ Acknowledgment
We thank Charles London, Zohar Ringel, and Shuofeng Zhang for their helpful comments. NF acknowledges the UKRI support through the Horizon Europe guarantee Marie Skłodowska-Curie grant (EP/X036820/1). SL was supported by the National Research Foundation of Korea (NRF) grants No.2020R1A5A1016126 and No.RS–2024-00462910. CM acknowledges funding from an EPSRC iCASE grant with IBM (grant number EP/S513842/1). SH gratefully acknowledges partial support from NSF grants DMS-2209975, 2015341, 20241842, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, and NSF grant MC2378 to the Institute for Artificial CyberThreat Intelligence and OperatioN (ACTION). The authors would also acknowledge support from His Majesty’s Government in the development of this research.
@@ Appendix A Glossary
$C$ Class count The number of classes. Section˜2
$p$ Layer width The width of the last layer Section˜2
$n$ Sample count The number of training samples Section˜2
$\mathcal{X}$ Input space Space of inputs Section˜2
$q$ Input distribution The probability distribution that generates samples in the input space. Section˜2
$\Phi$ Feature map A map from input to post-activation of the penultimate layer (the activations fed to the last layer). Section˜2
@@
A $p$-dimensional random variable or the post-activation of penultimate layer for $x\sim q$. Section˜2
$\mathcal{T}$ (Feature kernel) integral operator A map from Hilbert space to Hilbert space that depends on $\Phi(x),~x\sim q$ (typically the last layer features) Eq.˜1
$\mathcal{T}_{MP}$ Minimum projection operator A special set of operators that depends on the learned function $\hat{f}$. Definition˜1
$e_{k}$ $k^{\mathrm{th}}$ eigenfunction The eigenfunction of $\mathcal{T}.$ Eq.˜2
$\rho_{k}$ $k^{\mathrm{th}}$ eigenvalue The eigenvalue of $\mathcal{T}$. Eq.˜2
$\hat{f}$ Learned function A $C$-dimensional vector output function $\hat{f}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ expressed by the neural network. Section˜2
$f^{*}$ Target function A $C$-dimensional vector output function $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{C}$ with correct labels. The output is always a one-hot vector (up to a scaling constant). Section˜2
$\mathcal{H}^{*}$ Target function space Space linearly spanned by the entries of the target function. Section˜3
$\hat{\mathcal{H}}$ Learned function space Space linearly spanned by the entries of the learned function Section˜3
$P_{\mathcal{H}}$ Projection operator Projection operator onto $\mathcal{H}$. If $\mathcal{H}=span\{e_{1},e_{2},\ldots,e_{p}\}$, where $e_{k}$’s are orthonormal, the projection operator is given as $|P_{\mathcal{H}}=\sum_{k=1}^{p}\ket{e_{k}}\bra{e_{k}}$. Section˜3
$\alpha$ Downscale constant Target downscaling prefactor: $y\rightarrow y/\alpha$ Section˜4.1
$\Sigma_{b}$ Inter-class covariance matrix Covariance of class mean vectors for the training set. Eq.˜34
$\Sigma_{W}$ Intra-class covariance matrix Covariance of training feature vectors within given class. Eq.˜33
@@
$\|\theta\|_{F}^{2}$ Parameter norm Norm of all parameters in the model. It was used as richness measure with smaller value meaning richer dynamics. Section˜4.1
$\mathcal{D}_{LR}$ low rank metric Our proposed metric of dynamical richness. Smaller is richer. Eq.˜4
$\mathcal{S}_{init}$ Kernel deviation The CKA measure between the learned and initial kernel. Section˜4.1
$NC1$ Neural collapse metric The trace of $\Sigma_{b}^{\dagger}\Sigma_{W}$, measuring the training feature vectors variance compared to class boundaries. Proposition˜1
$\Pi^{*}(k)$ Cumulative quality Measure of how well the first $k$ eigenfunctions span the target function space. Eq.˜5
$\hat{\Pi}(k)$ Cumulative utilization Measure of how well the first $k$ eigenfunctions span the learned function space. Eq.˜5
$Q^{*}(k)$ Quality Per-feature quality or $\Pi^{*}(k)-\Pi^{*}(k-1)$ where $\Pi^{*}(0)=0$ Fig.˜5
$\hat{Q}(k)$ Utilization Per-feature utilization or $\hat{\Pi}(k)-\hat{\Pi}(k-1)$ where $\hat{\Pi}(0)=0$ Fig.˜5
@@ B.1 Bra-ket notation
$\mathcal{T}[f](x)=[\Phi_{1}(x)\braket{\Phi_{1}|f},\ldots,\Phi_{p}(x)\braket{\Phi_{p}|f}],$
where $\braket{\Phi_{k}|f}$ are scalars and $\Phi_{k}:\mathcal{X}\rightarrow\mathbb{R}$ are functions.
@@ B.2 Features are random variables and (well-behaved) random variables form a Hilbert space
The distinction becomes more important when we wish to discuss Hilbert space. A Hilbert space requires an inner product between functions (feature maps), which depends on the underlying distribution. For example, $\sin(x)$ and $\sin(2x)$ are orthogonal for $q=unif[0,2\pi]$ where $\int_{0}^{2\pi}\sin(x)\sin(2x)dx=0$, but not for unit Gaussian distribution $q=\mathcal{N}(0,1)$ where $\int_{-\infty}^{\infty}\sin(x)\sin(2x)e^{-x^{2}/2}dx\neq 0$.
For an underlying probability distribution and a set of functions (feature maps), we can form a Hilbert space. A common example of a Hilbert space is the set of solutions expressed by a linear model such as
$f(x)=w_{1}x+w_{2}x^{2}+w_{3}x^{3}.$
Assuming the features $x,x^{2},x^{3}]$ are linearly independent on input distribution $q$, they form a 3-dimensional Hilbert space.
B.3 Centered kernel alignment
The centered kernel alignment (CKACKA(Kornblith et al., 2019) measures the similarity between two matrices or operators, and are commonly used for tracking the evolution of NTK (Baratin et al., 2021; Lou et al., 2022):
$CKA(A,B)=\frac{\mathrm{Tr}(c(A)c(B))}{\|c(A)\|_{F}\|c(B)\|_{F}},$ (7)
@@ B.3 Centered kernel alignment
For $\mathcal{T}_{MP}$, if the learned function space $\hat{\mathcal{H}}$ contains the constant function — which is empirically true throughout the training for all our experiments and always true if it perfectly fits the training samples (Appendix˜D) — $c(\mathcal{T}_{MP})/\|c(\mathcal{T}_{MP})\|_{F}$ becomes an orthogonal projection operator on the constant complement of $\hat{\mathcal{H}}$, and is independent of the values of $a_{1}$ and $a_{2}$.
@@ B.4 Effective dimension
A covariance matrix or an integral operator may be fully ranked, but their eigenvalues may decay fast (e.g., by a power-law). The small eigendirections are often negligible, and researchers use various effective dimensions to measure the number of significant dimensions. In this appendix, we will use the exponent of entropy (Hill, 1973) to measure the effective dimension.
For a matrix or an operator with positive eigenvalues $\rho_{1},\cdots,\rho_{p}]$, the effective dimension is
@@
The effective dimension of a matrix (operator) can be interpreted as the number of linearly independent vectors (functions) needed to effectively describe the matrix (operator). This is similar in spirit to the principal component analysis (PCA) (Abdi and Williams, 2010) in that only the directions with significant variance are considered.
As discussed in Appendix˜D, the first eigenfunction is always a constant function, which is irrelevant for our measure $\mathcal{D}_{LR}$ and dynamical interpretation. Because the first eigenvalue is often much larger than other significant eigenvalues, we use
@@ Appendix C Linear models, features, and their inductive bias
Here, we will introduce the recent findings on the inductive bias of linear models and demonstrate the significance of eigenvalues and eigenfunctions of $\mathcal{T}$. Linear regression or kernel regression is a rare case where its inductive bias is analytically calculable (Bordelon et al., 2020; Jacot et al., 2020; Spigler et al., 2020; Canatar et al., 2021; Cui et al., 2021; El Harzli et al., 2024; Simon et al., 2023). For a given feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$ or kernel $K(x,x^{\prime})=\Phi(x)^{T}\Phi(x^{\prime})$, the model is expressed as the $p$ dimensional ($p$ can be infitnite) linear model
@@ C.1 Eigenfunctions - orthonormalized features
In many cases, $\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)]$ are not orthonormal. For example, the features $x,x^{2}]$ of $f(x)=w_{1}x+w_{2}x^{2}$ for $q=\text{Uniform}[0,1]$ are not orthonormal:
$\braket{x|x^{2}}=\int_{0}^{1}x^{3}dx\neq 0,\quad\braket{x|x}=\int_{0}^{1}x^{2}dx\neq 1,\quad\braket{x^{2}|x^{2}}=\int_{0}^{1}x^{4}dx\neq 1.$ (12)
The diagonalization into the eigenfunctions returns the orthonormal basis of the hypothesis (Hilbert) space $\mathcal{H}=\mathrm{span}\{\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)\}$, more formally known as Reproducing Kernel Hilbert Space (RKHS). The eigenvalue now shows the norm of the feature along the direction of eigenfunction as the features $\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)]$ are not normalized (Fig.˜6).
Figure 6: Diagonalization of features Since the features $\Phi_{1}(x),\Phi_{2}(x),\cdots,\Phi_{p}(x)]$ span a vector space, each function $\Phi_{k}:\mathcal{X}\rightarrow\mathbb{R}$ can be represented as a vector. The overlap among $\Phi_{k}$’s (non-zero linear correlation) and differing norms of $\Phi_{k}$’s result in diagonalized features $\sqrt{\rho_{1}}e_{1},\sqrt{\rho_{2}}e_{2},\cdots,\sqrt{\rho_{p}}e_{p}]$ to have distinct intensities (norms) $\rho_{1},\cdots,\rho_{p}$.
@@
$\Phi_{i}(x)=\sum_{j}O_{ij}\sqrt{\rho_{j}}e_{j}(x),\quad\quad O_{ij}=\frac{1}{\sqrt{\rho_{j}}}\braket{\Phi_{i}|e_{j}},$ (13)
where $O\in\mathbb{R}^{p\times p}$ is an orthogonal matrix (follows trivially from Eq.˜2). Using Eq.˜13, the model can be reparameterized in the eigenfunction basis
@@
Note that because $\theta=Ow$, the norm of the parameters is conserved (i.e. $\|\theta\|_{F}=\|w\|_{F}$), indicating that weighted basis $\sqrt{\rho_{1}}e_{1},\sqrt{\rho_{2}}e_{2},\cdots,\sqrt{\rho_{p}}e_{p}]$ instead of normalized basis $e_{1},e_{2},\cdots,e_{p}]$ must be used to represent the ‘intensity’ or norm of the features $\Phi_{1}(x),\cdots,\Phi_{p}(x)]$.